OpenStudy (anonymous):

Solve for x: 5x + 2 = 4x − 9

OpenStudy (anonymous):

5x-4x = -9-2 x = -11

OpenStudy (anonymous):

5x-4x = -9-2 x = -11

OpenStudy (anonymous):

tahnkyou :D